# Recognition-Synergistic Scene Text Editing

### v1

@@ 1 Introduction
  • We propose a simple yet effective scene text editing method dubbed RS-STE, which conducts conducts recognition-synergistic scene text editing in a unified framework. Such design enables implicit separation between background style and text content, thereby eliminating intricate model design.
@@ 2.1 Scene Text Editing
Recently, several works  [16, 4, 5, 39] explore the manipulation of text within the entire image based on stable diffusion [33]. They encompass the intricate aspects of layout arrangement and the accurate rendering of textual content. As this focus differs from the concerns addressed in our paper, we have not conducted a comparison with our work.
@@ 3.2 RS-STE
where $\mathbf{D}^{\text{t}}_{T_{A}}\in\mathbb{R}^{L\times C}$ and $\mathbf{D}^{\text{i}}_{I_{B}}\in\mathbb{R}^{N\times C}$ are the decoded token features for $T_{A}$ and $I_{B}$, respectively. It is noteworthy that the first $(L+N)$ output tokens aligned with $\mathbf{E}^{\text{t}}_{T_{B}}$ and $\mathbf{E}^{\text{i}}_{I_{A}}$, denoted as $\mathbf{D}_{\text{NULL}}$, are not used since they cannot access the full content of $\mathbf{E}^{\text{t}}_{T_{B}}$ and $\mathbf{E}^{\text{i}}_{I_{A}}$. $\mathbf{D}^{\text{t}}_{T_{A}}$ is further used to perform text recognition by a fully connected layer (FCN) that predicts the character probabilities $\mathbf{P}_{T_{A}}\in\mathbb{R}^{L\times{(|\Sigma|+1})}$, while $\mathbf{D}^{\text{i}}_{I_{B}}$ is fed into the Image Detokenizer to synthesize the edited image $I_{B}$.
@@ 4.3 Evaluation Metrics
Regarding the paired evaluation datasets Tamper-Syn2k and ScenePair, where the target edited images are specified, we utilize four metrics to assess the differences between the predicted and target images. These metrics include 1) Mean Squared Error (MSE), which measures the $L2$ distance; 2) Peak Signal-to-Noise Ratio (PSNR), representing the ratio of peak signal power to noise power; 3) Structural Similarity Index Measure (SSIM), which evaluates mean structural similarity; and 4) Fréchet Inception Distance [14] (FID), which calculates the distance between features extracted using the InceptionV3 model. Higher PSNR, SSIM, and lower MSE, FID indicate better performance.
@@ 4.5 Ablation Study
In the cyclic fine-tuning stage, cyclic perceptual loss ($\mathcal{L}^{\text{cyc}}_{\text{per}}$) and cyclic MSE loss ($\mathcal{L}^{\text{cyc}}_{\text{mse}}$) supervise the pixels of the reproduced results. Without these losses, as shown in Table 4 w/o $\mathcal{L}^{\text{cyc}}_{\text{per}}$ and w/o $\mathcal{L}^{\text{cyc}}_{\text{mse}}$, though RS-STE can achieve a better recognition accuracy, it fails to preserve the style and tend to generate the targeted content in an extraneous style, as illustrated in Figure 5 (b) w/o $\mathcal{L}^{\text{cyc}}_{\text{per}}$.
@@ Appendix A Summary
This supplementary material comprises four components: (1) detailed descriptions of MMPD in our RS-STE; (2) implementation details of the fine-tuning stage of detokenizer and data augmentation for recognition; (3) additional ablation studies on model size and the feature representation approach; and (4) more visualization examples generated by various scene text editing methods and our RS-STE.
@@ Appendix B Details of MMDP
As described in Section  3, the input of MMPD can be denoted as $[\mathbf{E}^{\text{t}}_{T_{B}},\mathbf{E}^{\text{i}}_{I_{A}},\mathbf{E}^{\text{t}}_{\text{query}},\mathbf{E}^{\text{i}}_{\text{query}}]\in\mathbb{R}^{2(L+N)\times C}$, where $L$ presents the length of the text embeddings and $N$ presents the length of the flattened image embeddings. In our configuration, we set $L=32$, $N=256$ and $C=384$. Our MMPD consists of 12 transformer blocks, each of which includes a layer normalization layer, causal self-attention with 6 heads, and a fully connected layer.
@@ Appendix C More Implementation Details
Fine-tuning Stage of Detokenizer. We initialize the pre-trained VAE from LDM [33] using configuration parameters $f=4$, $Z=8192$ and $d=3$. To improve the decoder’s performance in reconstructing text images from continuous features, we fine-tune the VAE on our training datasets. Specifically, we remove the codebook-related components from the pre-trained model and train it for 100k iterations using the Adam [kingma2014adam] optimizer with a batch size of 256, and a learning rate of $1.25\times 10^{-3}$. The reconstruction performance of the VAE before and after fine-tuning on the evaluation dataset Tamper-Syn2k and ScenePair is shown in Tables 6 and  7. Compared to the pre-trained VAE, the fine-tuned VAE demonstrates better image reconstruction performance for text images. This metric also indicates the upper limit of the image editing performance when using the VAE decoder as an Image Detokenizer.
@@ D.2 Discrete Feature Representation
This can primarily be attributed to two factors: (1) The discretization of images introduces information distortion, resulting in poorer reconstruction quality compared to continuous representations. As shown in Table 9, for the given dataset, the reconstruction performance of the discrete form is inferior to that of the continuous form. (2) Continuous representations effectively mitigate the inherent decoding bias of the detokenizer. As discussed in Section  3.3, for continuous image features, reconstruction loss can be computed on the detokenized images, ensuring pixel-level accuracy in the final output. In contrast, for discrete representations, supervision can only be applied to the discretized image features decoded by the MMPD, leading to feature distortions during the detokenization process.
@@ D.3 Loss Weights
In the cyclic training stage described in Section  4, we observe that the ratio of the recognition loss weight, defined as $\lambda_{\text{rec}}=(\lambda_{6}+\lambda_{7})/2$, to the image reconstruction loss weight, defined as $\lambda_{\text{recon}}=(\lambda_{4}+\lambda_{5})/2$, plays a crucial role in ensuring content and style consistency. Consequently, we conduct an ablation study to examine the effects of varying this ratio, as shown in Figure 7. Our findings indicate that a ratio close to 10 consistently produces high-quality images.
@@ Appendix E Limitation and Analysis
A potential limitation of our method as well as most other methods for scene text editing lies in the limited performance when editing images with extremely large text curvature, as shown in Fig  9. This limitation is mainly attributed to the scarcity of such data in synthetic training data. To further investigate this issue, we train our model with additionally synthetic curved text samples generated using the synthesis engine mentioned in line 338, and our method (RS-STE+) achieves robust curved text editing, which implies that such limitation arises from insufficient training data of curved text.

### v2

@@ 1 Introduction
  • We propose a simple yet effective scene text editing method dubbed RS-STE, which conducts recognition-synergistic scene text editing in a unified framework. Such design enables implicit separation between background style and text content, thereby eliminating intricate model design.
@@ 2.1 Scene Text Editing
Recently, several works  [16, 4, 5, 41] explore the manipulation of text within the entire image using stable diffusion [34]. They encompass the intricate aspects of layout arrangement and the accurate rendering of textual content. As this focus differs from the concerns addressed in our paper, we have not conducted a comparison with our work.
@@ 3.2 RS-STE
where $\mathbf{D}^{\text{t}}_{T_{A}}\in\mathbb{R}^{L\times C}$ and $\mathbf{D}^{\text{i}}_{I_{B}}\in\mathbb{R}^{N\times C}$ are the decoded token features for $T_{A}$ and $I_{B}$, respectively. It is noteworthy that the first $(L+N)$ output tokens aligned with $\mathbf{E}^{\text{t}}_{T_{B}}$ and $\mathbf{E}^{\text{i}}_{I_{A}}$, denoted as $\mathbf{D}^{\text{t}}_{\text{NULL}}$ and $\mathbf{D}^{\text{i}}_{\text{NULL}}$, are not used since they cannot access the full content of $\mathbf{E}^{\text{t}}_{T_{B}}$ and $\mathbf{E}^{\text{i}}_{I_{A}}$. $\mathbf{D}^{\text{t}}_{T_{A}}$ is further used to perform text recognition by a fully connected layer (FCN) that predicts the character probabilities $\mathbf{P}_{T_{A}}\in\mathbb{R}^{L\times{(|\Sigma|+1})}$, while $\mathbf{D}^{\text{i}}_{I_{B}}$ is fed into the Image Detokenizer to synthesize the edited image $I_{B}$.
@@ 4.3 Evaluation Metrics
Regarding the paired evaluation datasets Tamper-Syn2k and ScenePair, where the target edited images are specified, we utilize four metrics to assess the differences between the predicted and target images. These metrics include 1) Mean Squared Error (MSE), which measures the $L2$ distance; 2) Peak Signal-to-Noise Ratio (PSNR), representing the ratio of peak signal power to noise power; 3) Structural Similarity Index Measure (SSIM), which evaluates mean structural similarity; and 4) Fréchet Inception Distance [14] (FID), which calculates the distance between features extracted using the InceptionV3 [39] model. Higher PSNR, SSIM, and lower MSE, FID indicate better performance.
@@ 4.5 Ablation Study
In the cyclic fine-tuning stage, cyclic perceptual loss ($\mathcal{L}^{\text{cyc}}_{\text{per}}$) and cyclic MSE loss ($\mathcal{L}^{\text{cyc}}_{\text{mse}}$) supervise the pixels of the reproduced results. Without these losses, as shown in Table 4 w/o $\mathcal{L}^{\text{cyc}}_{\text{per}}$ and w/o $\mathcal{L}^{\text{cyc}}_{\text{mse}}$, though RS-STE can achieve a better recognition accuracy on Tamper-Scene, it fails to preserve the style and tend to generate the targeted content in an extraneous style, as illustrated in Figure 5 (b) w/o $\mathcal{L}^{\text{cyc}}_{\text{per}}$.
@@ Appendix A Summary
This supplementary material comprises four components: (1) detailed descriptions of MMPD in our RS-STE; (2) implementation details of the fine-tuning stage of detokenizer and data augmentation for recognition; (3) additional ablation studies on model size and the feature representation approach; (4) limitation and analysis; and (5) more visualization examples generated by various scene text editing methods and our RS-STE.
@@ Appendix B Details of MMDP
As described in Section 3, the input of MMPD can be denoted as $[\mathbf{E}^{\text{t}}_{T_{B}},\mathbf{E}^{\text{i}}_{I_{A}},\mathbf{E}^{\text{t}}_{\text{query}},\mathbf{E}^{\text{i}}_{\text{query}}]\in\mathbb{R}^{2(L+N)\times C}$, where $L$ presents the length of the text embeddings and $N$ presents the length of the flattened image embeddings. In our configuration, we set $L=32$, $N=256$ and $C=384$. Our MMPD consists of 12 transformer blocks, each of which includes a layer normalization layer, causal self-attention with 6 heads, and a fully connected layer.
@@ Appendix C More Implementation Details
Fine-tuning Stage of Detokenizer. We initialize the pre-trained VAE from LDM [34] using configuration parameters $f=4$, $Z=8192$ and $d=3$. To improve the decoder’s performance in reconstructing text images from continuous features, we fine-tune the VAE on our training datasets. Specifically, we remove the codebook-related components from the pre-trained model and train it for 100k iterations using the Adam [22] optimizer with a batch size of 256, and a learning rate of $1.25\times 10^{-3}$. The reconstruction performance of the VAE before and after fine-tuning on the evaluation dataset Tamper-Syn2k and ScenePair is shown in Tables 6 and 7. Compared to the pre-trained VAE, the fine-tuned VAE demonstrates better image reconstruction performance for text images. This metric also indicates the upper limit of the image editing performance when using the VAE decoder as an Image Detokenizer.
@@ D.2 Discrete Feature Representation
This can primarily be attributed to two factors: (1) The discretization of images introduces information distortion, resulting in poorer reconstruction quality compared to continuous representations. As shown in Table 9, for the given dataset, the reconstruction performance of the discrete form is inferior to that of the continuous form. (2) Continuous representations effectively mitigate the inherent decoding bias of the detokenizer. As discussed in Section 3.3, for continuous image features, reconstruction loss can be computed on the detokenized images, ensuring pixel-level accuracy in the final output. In contrast, for discrete representations, supervision can only be applied to the discretized image features decoded by the MMPD, leading to feature distortions during the detokenization process.
@@ D.3 Loss Weights
In the cyclic training stage described in Section 3.3, we observe that the ratio of the recognition loss weight, defined as $\lambda_{\text{rec}}=(\lambda_{6}+\lambda_{7})/2$, to the image reconstruction loss weight, defined as $\lambda_{\text{recon}}=(\lambda_{4}+\lambda_{5})/2$, plays a crucial role in ensuring content and style consistency. Consequently, we conduct an ablation study to examine the effects of varying this ratio, as shown in Figure 7. Our findings indicate that a ratio close to 10 consistently produces high-quality images.
@@ Appendix E Limitation and Analysis
A potential limitation of our method as well as most other methods for scene text editing lies in the limited performance when editing images with extremely large text curvature, as shown in Figure 9. This limitation is mainly attributed to the scarcity of such data in synthetic training data. To further investigate this issue, we train our model with additionally synthetic curved text samples generated using the synthesis engine mentioned in Section 4.1, and our method (RS-STE+) achieves robust curved text editing, which implies that such limitation arises from insufficient training data of curved text.